3) Serena Company has budgeted the following costs for the production of its only product:

Direct Materials $35,000

Direct Labor 25,000

Variable indirect production costs 30,000

Fixed indirect production costs 15,000

Variable selling and administrative costs 7,500

Fixed selling and administrative costs 12,500

Total Costs $125,000

Revenue is $200,000

Serena Company has a target profit of $50,000. What is the average target markup percentage for setting prices as a percentage of variable manufacturing costs? calculate the operating income using contribution and absorption methods

A) 69%

B) 158%

C) 24%

D) 94%
Answer #1

Direct Materials= $35,000

Direct Labor = $25,000

Variable indirect production costs = $30,000

Variable manufacturing cost = Direct Materials+Direct Labor+Variable indirect production costs

= 35,000+25,000+30,000

= $90,000

Fixed indirect production costs = $15,000

Variable selling and administrative costs = $7,500

Fixed selling and administrative costs = $12,500

Target Profit = $50,000

Target markup percentage = ( Fixed indirect production costs+Variable selling and administrative costs+Fixed selling and administrative costs +Target Profit) / Variable manufacturing cost

= ( 15,000+7,500+12,500+50,000)/90,000

= 85,000/90,000

= 94%

Correct option is D.

Kindly give a positive rating if you are satisfied with this solution and please ask if you have any query.

Thanks